what are x and y in the following two peoblems?
6xt40y=554 and 8x+20y=472

Answer by rothauserc(4718)

1) 6x +40y = 554
:
2) 8x+ 20y = 472
:
multiply equation 2 by 2
:
3) 16x +40y = 944
:
subtract equation 1) from equation 3)
:
4) 10x = 390
:
x = 39
:
substitute for x in equation 1)
:
6(39) +40y = 554
:
234 +40y = 554
:
40y = 320
:
y = 8
:
*************************************************************
x = 39 and y = 8
:
check the solution by substituting in equation 2 for x and y
:
8(39) + 20(8) = 472
:
312 + 160 = 472
:
472 = 472
:
our answer checks
**************************************************************
: